OpenStudy (anonymous):

can the following numbers \(23,45326,56740,16320\) be written as a sum of two squares,explain!!!

OpenStudy (anonymous):

this are 4 seperate numbers ,NOT one number

OpenStudy (whpalmer4):

At least one of them can be. Are there any restrictions on what we are squaring?

OpenStudy (whpalmer4):

For example, you can write 23 as \[23=(\frac{5}{2})^2 + (\frac{\sqrt{67}}{2})^2\] if you allow the numbers to be any reals...

OpenStudy (whpalmer4):

if you only allow integers, however, then the answer is that 23 cannot be written as the sum of two squares.

OpenStudy (anonymous):

we are squaring integers here ,forgot to mention that

OpenStudy (anonymous):

so the first thing i see in most questions ,is testing wether the number is Gaussian prime, \(n\equiv 2,3 \mod 4\) gaussian prime ,\( n\equiv 1 \mod 4\) not gaussian prime

OpenStudy (anonymous):

actually for primes ,not just any \(n\),also the condition is the other way around,meaning if\(p\equiv 1\mod 4 \) then \(p\) is not just an ordianry prime but gaussian prime,hence the prime can be written as a sum of two squares,i am interested in factorising this numbers